OpenStudy (anonymous):

The sum of two numbers is 56. Twice the smaller number exceeds one half the larger number by 22. Find the numbers.

OpenStudy (anonymous):

do you know how to write that algebraically?

OpenStudy (anonymous):

Wow I'm probably very wrong :/ x + y =56

OpenStudy (anonymous):

that's one equation

OpenStudy (anonymous):

solve the following two linear equations let the larger number is X and smaller is Y X+Y=56 2Y-0.5X=22

OpenStudy (anonymous):

or 2y = 1/2 x + 22 same shiz

OpenStudy (anonymous):

Thanks! I'll take it from there.